OpenStudy (anonymous):

f ' (x) = 2(2xsin(x)+x^2cos(x)) f ' (3)=?? Help please!

OpenStudy (anonymous):

i think you have to plug in the 3 in the x's of the function... so you set it up like f(3)=??????... yes... then solve and continue...

OpenStudy (anonymous):

Yeah, but what about sin(3) and cos(3), there's no such thing.

hero (hero):

f ' (x) = 2(2xsin(x)+x^2cos(x)) f'(3) = 2(2(3)sin(3)+3^2cos(3)) f'(3) =

hero (hero):

There is a such thing sin(3) and cos(3)

OpenStudy (anonymous):

oops... Thanks, I Hate MATH, too. Haha